# Turnstile area as a measure for chaotic transport in magnetic confinement fusion devices

### v1

@@ VI.1 Turnstile area in fields from the QUASR database
The QUASR database contains roughly 300,000 quasi-axisymmetric stellarator configurations [30, 31]. These configurations each consist of a set of coils and currents that have been optimized to produce a vacuum magnetic field that has been optimized to confine particles through quasisymmetry [46, 47]. The entire database can be browsed interactively https://quasr.flatironinstitute.org/, and contains configurations with different numbers of coils, field periods and types of symmetry. The coils can be downloaded in the format used by simsopt.
We focus on one specific QUASR configuration with identifier 242612, which is a four field period device with only eight coils in total. The rotational transform starts below 1, and monotonically increases to a value of around 1 at the edge. We plot the coils and the field strength a surface of the configuration in Figure 7 a-b. The coils are indicated with gray curves, and a magnetic surface of the configuration is shown, colored by the magnitude of the magnetic field on the surface.
Figure 7: Analysis of QUASR configuration 242612. a, b: Coils and a magnetic surface of the configuration colored by the magnetic field strength. c: Poincaré section at $\phi=0$ plane, including the manifolds of the $9/10$ island chain and the outer $/4$ 4 island chain, as well as manifolds around the hyperbolic point inside of the $/4$ 4 island chain. d: zoom in of the turnstiles around the edge $/4$ 4 island. d: rotational transform profile calculated along the midplane.
The Poincaré section of the field, shown in figure 7 (c), shows a few islands in the core, and a large $/4$ 4 island chain at the edge, which is surrounded by a stochastic region. This stellarator configuration thus resembles a conventional stellarator with an island divertor like W7-X, except optimized for quasisymmetry instead of quasi-omnigeneity. One interesting feature of this configuration is that the resonance in the edge has itself bifurcated, resulting in two elliptic fixed points shown in light and dark green, and a hyperbolic point in between.
We calculate the manifolds, heteroclinic points, and turnstiles for both the inner and outer heteroclinic connections in the 9/10 and 4/4 resonances, as well as the period four homoclinic connections (left and right). The results of this are presented in table 1
@@
xp9manifold 6.3879e-12 3.7425e-12
outermanifold 9.1613e-07 3.4471e-05
islandmanifold 1.3626e-06 1.3627e-06
@@ VI.2 Turnstile area calculation verification
We next use a more robust method, where we actually resolve the turnstile lobe using 40 trajectories logarithmically spaced between the heteroclinic orbits. We calculate the area using Gauss’s area formula (also known as the shoelace algorithm) which gives the area enclosed by a polygon determined by points ${x_{i},y_{i}}$ by:
@@
This robust method is also used in the code lober [15] to evaluate lobe areas. The result of this calculation is shown in table 2.
@@
The last verification that we do is to directly calculate $\oint_{\rm lobe}\mathbf{A}\cdot\mathrm{d}l$ over the boundary of the lobe by a simple Riemannian sum over the segments of the lobes.
The four different calculations show good correspondence, with the crude triangle approximation of course being the least accurate.
@@ VI.3 QUASR 229079
We next analyze the structures in the edge of QUASR configuration 229079. This is a three field-period quasi axisymmetric stellarator with 18 coils and a low rotational transform. The coils and a flux surface of this configuration are shown in figure 9. This configuration is rather compact, and has strongly shaped ridges on the inboard side of the tori which create rotational transform, very similar to the configurations generated by Plunk [48] through a perturbative modification of axisymmetric equilibria. Configurations such as this, though not explicitly optimized for, are common in the QUASR database among the high aspect-ratio low rotational transform configurations. Henneberg and Plunk [49] have investigated the potential of such configurations as a reactor dubbed the ’Compact Stellarator-Tokamak hybrid’.
@@
Figure 9: Analysis of QUASR configuration 229097. a, b: Coils and a magnetic surface of the configuration colored by the magnetic field strength. c: Poincaré section at $\phi=0$ plane, showing an ordered center surrounded by a large stochastic region that contains many high-order resonances. The manifolds associated with the 6/16 (light blue/blue and violet/purple), 3/9 (light-/dark green), 3/10 (yellow/orange) , and 3/11 (pink/fuchsia) are shown, as well as the fixed points of a 3/12 chain in pink. c: zoom in of the interacting turnstiles in the chaotic edge, with only the manifolds around a single island resonance zone colored. d: rotational transform profile calculated along the midplane.
In the Poincaré plot of this configuration we see, towards the edge a few islands, starting with the 3/7 and the 6/15, and then the first resonance for which we analyze the turnstiles is the resonance on the 3/8 island, which is period-doubled and shows a 6/16 island chain. In such a case, the rational surface has broken into two separate islands, and we plot the stable (unstable) manifolds surrounding the first island with lavender (purple), the stable (unstable) manifolds around the second with light blue (blue). In such a period-doubled case, the turnstiles around one island chain can be much larger than those around the other.
The 6/16 island chain marks one of the last occurrences of clear surfaces surrounding the plasma, what follows is a very large region of stochastic magnetic field. In this region many resonances can be found, of which we show the 3/9, 3/10, 3/11 in different colors.
@@
Figure 7 d shows a zoom in to these manifolds, with only one occurrence of each colored. We see the size of the manifold with respect to the size of the resonance zone increase significantly from each resonance zone to the next. This is quantified in table 3. The size of the island even becomes significant with respect to the size of the resonance zone, and we even see a large fraction of overlap between the manifolds of the 3/10 and 3/11 resonances.
@@
manifold8a $075.8090e-$ 07 $062.5502e-$ 06
manifold8b $105.2863e-$ 10 $104.4831e-$ 10
manifold9 $053.6184e-$ 05 $057.3541e-$ 05
manifold10 $041.3430e-$ 04 $042.1494e-$ 04
manifold11 $042.5547e-$ 04 $043.7709e-$ 04
@@
The rotational transform profile is calculated by integrating over 40 field periods, and fitting the rotation angle. This is shown in figure 9 d). This is not a perfect measure for rotational transform, but the profile does show that the configuration crosses the 1/4 resonance twice, and in the Poincaré plot we do find two 3/12 resonances the outer one is shown with pink markers at the very edge of the configuration. This indicates that there are nontwist map dynamics [26] at play in this outermost stochastic region.
@@ VI.4 Turnstile area in the chaotic low-iota configurations in the W7-X stellarator
Finally we will use the tools to calculate the turnstile area in configurations of the Wendelstein 7-X (W7-X) stellarator. W7-X was optimized for neoclassical transport and low bootstrap current. It is therefore said to be a quasi-isodynamic stellarator. W7-X has already shattered the records for stellarator field accuracy [50], fusion triple product [51], and confinement [52].
W7-X is a five field period device that creates the magnetic field with a total of 70 main coils222there are also five windowpane coils and ten trim coils that can be used to further adjust the field, but do not generate the bulk of the confinement field that generate the field. Per half-field period, there are 5 unique modular coils, that have complicated shapes and two planar coils. The coil configuration and the shape for the plasma region is shown in figure 10 b. These are copied around the device according to stellarator symmetry and field periodicity to produce the full coil set. The power supplies to these superconducting coils are connected according to this symmetry, but the current in each of these 7 unique coil groups can be set separately.
When the modular coils carry their design current of 1.62 MA each, and there is no current in the planar coils, the stellarator generates the standard configuration. This configuration has a large volume of closed flux surfaces, that is terminated with a large 5/5 island chain that acts as a divertor and determines the strikepoints of the plasma on the wall (and the location of heat loading).
@@
Figure 10: The W7-X GYM00+1750 configuration. a) Connection length plot showing the minimum distance (both forwards and backwards along field lines) to the wall. Overlaid are the unstable (stable) manifolds of the 5/4 island chain in white (orange). The features in the connection length plot correlate strongly with the manifolds of the outer heteroclinic connection. The inner heteroclinic connection has an small turnstile area compared to the outer heteroclinic connection. b: Modular coils (white) and planar coils(blue) of the W7-X stellarator, as well as a magnetic surface, colored by the magnetic field strength on it. c: Rotational transform in the GYM configuration. d: Poincaré section of the full $\phi=0$ cross section, showing nested surfaces inside and a very stochastic edge.
@@
One method that is often employed to design divertors is to analyze the connection length in the divertor region, with tools such as EMC3-EIRENE. The EMC3-Lite code is capable of calculating heat loads on plasma-facing components (PFCs) and PFC-to-PFC connection lengths for a given magnetic field and PFC arrangement [54]. For both, it relies on following Monte Carlo particles (which either simply follow magnetic field lines or take a random walk to imitate anisotropic diffusion), which are tracked on a field-aligned grid (for more details on the grid scheme see [55]). The PFC-to-PFC connection length $L_{c}$ is specific to each field line and is not recorded, but the grid cell-averaged connection length $<L_{c}>$ is calculated and recorded, and is useful for revealing the structure of the magnetic field and predicting regions of heat loading on PFCs (see for example [54, 56]). $<L_{c}>$ is calculated by tracing Monte Carlo particles both forwards and backwards along the field until they intersect a PFC. The connection length of this field line contributes to all grid cells it passes through, and $<L_{c}>$ for a given cell is the arithmetic mean of $L_{c}$ of all unique field lines passing through the cell. The cell in which each magnetic field line is initialised is selected to ensure that each cell contains at least $n_{\text{min}}$ unique trajectories. The location of an initialised field line in a cell is uniformly random. An upper cutoff $L_{c,\text{max}}$ is applied to each field line, since field lines which do not intersect a PFC (intact surfaces around island O-points, for example) would otherwise be traced indefinitely. In the result presented here we select $n_{\text{min}}=100$ and $L_{c,\text{max}}=10^{4}$ meters. The grid resolution used here is 480 points in the radial direction and 1025 points in the poloidal direction. This is several times larger than is typically required for heat deposition calculations but is used here to show the fine structure emerging from the magnetic chaos.
@@
The connection length in the island region is plotted in figure 10 a, and shows a remarkable pattern of alternating long connection length and short connection length regions. The manifolds of the outer heteroclinic connection are also shown on this plot, and there is a strong correlation between the two.
@@
The turnstile mechanism helps to explain how this pattern comes about: the region that is in neither lobe of the first many lobes of the heteroclinic tangle, but is close to the center of the island, has a high likely hood of remaining in that region, and thus has a large connection length. The region that is in one lobe, but not the other, has a probability of hitting the target, but only a fraction of the trajectories in each lobe does. This dark band of low connection length is thus mapped, over and over again, in an ever repeating pattern that Poincaré dubbed a ’trellis’ [10], made visible in our analysis of the stellarator divertor!
Just like the QUASR configurations, we again see that the turnstile area on the outer heteroclinic connection is much larger than the turnstile on the inner heteroclinic connection. The oscillations of the inner manifold are not visible at all at this scale of magnification. This means that transport into the island is likely governed by other mechanisms, such as turbulent transport and neoclassical effects.
Plasma that crosses this separatrix is channeled along the stable (unstable) manifold, where it can intersect the wall. The field is so designed that the divertor intersects this stream, and any heat that is transported through the inner turnstile is immediately transported to this wall, and the stochasticity originating from the turnstile on this outer heteroclinic connection is unlikely to have much effect on the heat transport to the wall.
@@
We have seen in all the stellarator configurations, that on each resonance the turnstile on the outer heteroclinic connection is significantly larger than the turnstile on the inner heteroclinic connection. This is likely because the perturbations that induce the stochasticity are in large part due to ripple effects caused by the finite number of discrete coils. The field lines forming the outer turnstile pass much closer to these coils, and are much more affected.
@@
In terms of stellarator divertor design, it would be interesting to understand if this feature is unavoidable, or just a feature of the configurations analyzed. In other words, would it be possible to generate a field with coils, where the turnstile on the inner separatrix is large, potentially larger than that of the outer separatrix. This could create a stochastic buffer region that could aid in achieving plasma detachment. Configurations such as the GYM00+1750, where the stochasticity is almost completely caused by the interactions of the outer manifolds, do not provide any significant change in the transport from plasma to wall.

### v3

@@ VI.1 Turnstile area in fields from the QUASR database
The QUASR database currently contains over 370,000 quasi-axisymmetric stellarator configurations [35, 36]. The entire database can be browsed interactively https://quasr.flatironinstitute.org/, and contains configurations with different numbers of coils, field periods and types of symmetry. Each configuration is defined by a set of coils (and their currents) that produce a quasisymmetric vacuum magnetic field, and a set of toroidal surfaces that were targeted in the optimization. The coils can be downloaded in the format used by simsopt. Though the agreement between the field and the surfaces was one of the targets in the optimization, this agreement is not always perfect, and many configurations do contain islands, especially if the rotational transform reaches a rational value on the outer surfaces, close to the coils.
We focus on one specific QUASR configuration with identifier 242612, which is a four field period device with only eight coils in total. The rotational transform starts below 1, and monotonically increases to a value of around 1 at the edge. We plot the coils and the field strength a surface of the configuration in Figure 7 (a)-(b). The coils are indicated with gray curves, and a magnetic surface of the configuration is shown, colored by the magnitude of the magnetic field on the surface.
[…]
Figure 7: Analysis of QUASR configuration 242612. (a), (b): Coils and a magnetic surface of the configuration colored by the magnetic field strength. (c): Poincaré section at $\phi=0$ plane, including the manifolds of the $9/10$ island chain and the outer $/4$ 4 island chain, as well as manifolds around the hyperbolic point inside of the $/4$ 4 island chain. (d): zoom in of the turnstiles around the edge $/4$ 4 island. (e): rotational transform profile calculated along the midplane.
The Poincaré section of the field, shown in figure 7 (c), shows a few islands in the core, and a large $/4$ 4 island chain at the edge, which is surrounded by a chaotic region. This stellarator configuration thus resembles a conventional stellarator with an island divertor like W7-X, except optimized for quasisymmetry instead of quasi-omnigeneity. One interesting feature of this configuration is that the resonance in the edge has itself bifurcated, resulting in two elliptic fixed points shown in light and dark green, and a hyperbolic point in between.
We calculate the manifolds, heteroclinic points, and turnstiles for both the inner and outer heteroclinic connections in the 9/10 and 4/4 resonances, as well as the period four homoclinic connections (since here there is no inner or outer, we label these left and right). The results of this are presented in table 1.
@@
manifolds 9/10 $6.3879\cdot 10^{-12}$ $3.7425\cdot 10^{-12}$
surrounding heteroclinic manifolds $9.1613\cdot 10^{-07}$ $3.4471\cdot 10^{-05}$
island homoclinic manifolds (left/right) $1.3626\cdot 10^{-06}$ $1.3627\cdot 10^{-06}$
@@ VI.2 Turnstile area calculation verification
We next use a more robust method, where we resolve the turnstile lobe using 40 trajectories between the heteroclinic orbits. We calculate the area using Gauss’s area formula (also known as the shoelace algorithm) which gives the area enclosed by a polygon determined by points ${x_{i},y_{i}}$ by:
@@
This method is also used in the code lober [32] to evaluate the area between manifolds. The result of this calculation is shown in table 2.
@@
The last verification is to directly calculate $\oint_{\rm lobe}\mathbf{A}\cdot\mathrm{d}l$ over the boundary of the lobe by a simple Riemannian sum over the segments of the lobes. The four different calculations show good correspondence, with the crude triangle approximation of course being the least accurate with an error of about 4%, with the Biot-Savart evaluation on the boundary having a relative difference of about $5\cdot 10^{-4}$ and the shoelace area formula a relative difference of $10^{-5}$.
@@ VI.3 QUASR 229079
We next analyze the structures in the edge of QUASR configuration 229079. This is a three field-period quasi axisymmetric stellarator with 18 coils and a low rotational transform. The coils and a flux surface of this configuration are shown in figure 9 (a)-(b). This configuration is rather compact, and has strongly shaped ridges on the inboard side of the tori which create rotational transform, very similar to the configurations generated by Plunk [62] through a perturbative modification of axisymmetric equilibria. Configurations such as this, though not explicitly optimized for, are common in the QUASR database among the high aspect-ratio low rotational transform configurations. Henneberg and Plunk [63] have investigated the potential of such configurations as a reactor dubbed the ’Compact Stellarator-tokamak hybrid’.
[…]
@@
Figure 9: Analysis of QUASR configuration 229097. (a), (b): Coils and a magnetic surface of the configuration colored by the magnetic field strength. (c): Poincaré section at $\phi=0$ plane, showing an ordered center surrounded by a large chaotic region that contains many high-order resonances. The manifolds associated with the 6/16 (light blue/blue and violet/purple), 3/9 (light-/dark green), 3/10 (yellow/orange) , and 3/11 (pink/fuchsia) are shown, as well as the fixed points of a 3/12 chain in pink. (d): zoom in of the interacting turnstiles in the chaotic edge, with only the manifolds around a single island resonance zone colored. (e): rotational transform profile calculated along the midplane.
In the Poincaré plot of this configuration we see closed nested surfaces in the center, with an occasional island, but further out the field becomes more and more chaotic. The innermost island chain for which we calculate manifolds, is a period-doubled 6/16 island chain, shown in blues and purples. After that there is little discernible structure in the Poincaré plot. Nevertheless, there are many fixed points in this region.
In this chaotic region, we can still calculate the approximate rotational transform profile. We do this by only integrating for 20 field periods, and finding the average angle a field line rotates around the axis for this number of mappings. This is shown in figure 9 (e). We also plot the rational numbers corresponding with specific rationals.
@@
Even though the field outside the 6/16 resonance is completely chaotic, we can still find hyperbolic fixed points associated with the crossed resonances, and we can calculate their manifolds and turnstile areas. Figure 9 (d) shows a zoom in to these manifolds, with only one occurrence of each colored. We see the turnstile area increase significantly with respect to the size of the resonance zone from each resonance zone to the next. This is quantified in table 3. We again observe that the turnstile area is larger on the outboard side of the configuration.
We can see that in the first island chains are clearly distinguishable, and separate, with a small turnstile area relative to the resonance zone area. In such a region, individual island chains act as partial barriers, and field lines must thread through two small turnstiles, both on the inboard and on the outboard side of the chain to pass through. As one moves further and further into the chaotic region, the manifolds of subsequent island chains start to align, they become large with respect to the resonance zone area. In this case there is a large exchange of flux between one side of the island and the other. We also see that the turnstile on one periodic orbit has a large overlap with the turnstile on the next periodic orbit. In this regime, the resonances do not act as barriers at all, and field lines can pass from one resonance zone to the next in a few mappings. We speculate that the onset of turnstile overlap between neighboring resonances can cause a change in the character of the transport.
It is also interesting to note that the rotational transform profile seems to increase towards the very edge of the region where the field line map can be calculated, and crosses the 3/12 resonance a second time. This indicates that there are nontwist map dynamics [45] at play in this outermost chaotic region. We can even locate the periodic orbit associated with the 3/12 resonance in this very edge of the configuration, but because this region is close to the coils, many of the field lines here do not make a full orbit, and the manifolds cannot be calculated.
@@
manifolds 6/15 a $5.8090\cdot 10^{-07}$ $2.5502\cdot 10^{-06}$
manifolds 6/15 b $5.2863\cdot 10^{-10}$ $4.4831\cdot 10^{-10}$
manifolds 3/9 $3.6184\cdot 10^{-05}$ $7.3541\cdot 10^{-05}$
manifolds 3/10 $1.3430\cdot 10^{-04}$ $2.1494\cdot 10^{-04}$
manifolds 3/11 $2.5547\cdot 10^{-04}$ $3.7709\cdot 10^{-04}$
@@ VI.4 Turnstile area in the chaotic low-iota configurations in the W7-X stellarator
Finally we will use the tools to calculate the turnstile area in configurations of the Wendelstein 7-X (W7-X) stellarator. W7-X was optimized for neoclassical transport and low bootstrap current. It is therefore said to be a quasi-isodynamic stellarator. W7-X has already shattered the records for stellarator field accuracy [64], stellarator fusion triple product [65], and stellarator confinement [21].
W7-X is a five field period device that creates the magnetic field with a total of 70 main coils222there are also five windowpane coils and ten trim coils that can be used to further adjust the field, but do not generate the bulk of the confinement field that generate the field. Per half-field period, there are 5 unique modular coils, that have complicated shapes and two planar coils. The coil configuration and the shape for the plasma region is shown in figure 10 (b). These are copied around the device according to stellarator symmetry and field periodicity to produce the full coil set. The power supplies to these superconducting coils are connected according to this symmetry, but the current in each of these 7 unique coil groups can be set separately.
When the modular coils carry their design current of 1.62 MA each, and there is no current in the planar coils, the stellarator generates the standard configuration. This configuration has a large volume of closed flux surfaces, that is terminated with a large 5/5 island chain that acts as a divertor and determines the strikepoints of the plasma on the PFCs (and the location of heat loading).
@@
When the current in the modular coils is set to $I_{\rm modular}=1.1095MA$, and the current in the planar coils is $I_{\rm planar}=-\tfrac{1}{3}I_{\rm modular}\simeq-0.3661MA$, we get a configuration where the rotational transform in the edge is $\imath=1.25$. A Poincaré section of this field is shown in figure 10 (d), and the resultant rotational transform profile is shown in figure 10 (c). This configuration is called GYM00-1750 and is studied in the context of nonresonant divertors. One of the interesting features of this configuration is that the edge of the field is chaotic, which has led to its study in the context of nonresonant divertors [26, 5].
@@
Figure 10: The W7-X GYM00+1750 configuration. (a): Connection length plot showing the minimum distance (both forwards and backwards along field lines) to the PFCs. Overlaid are the unstable (stable) manifolds of the 5/4 island chain in white (orange). The features in the connection length plot correlate strongly with the manifolds of the outer heteroclinic connection. The inner heteroclinic connection has an small turnstile area compared to the outer heteroclinic connection. (b): Modular coils (white) and planar coils (blue) of the W7-X stellarator, as well as a magnetic surface, colored by the magnetic field strength on it. (c): Rotational transform in the GYM configuration. (d): Poincaré section of the full $\phi=0$ cross section, showing nested surfaces inside and a very chaotic edge.
@@
One method that is often employed to design divertors is to analyze the connection length in the divertor region, with tools such as EMC3-Lite. The EMC3-Lite code is capable of calculating heat loads on plasma-facing components (PFCs) and PFC-to-PFC connection lengths for a given magnetic field and PFC arrangement [2]. For both, it relies on following Monte Carlo particles (which either simply follow magnetic field lines or take a random walk to imitate anisotropic diffusion), which are tracked on a field-aligned grid (for more details on the grid scheme see [67]). The PFC-to-PFC connection length $L_{c}$ is specific to each field line and is not recorded, but the grid cell-averaged connection length $<L_{c}>$ is calculated and recorded, and is useful for revealing the structure of the magnetic field and predicting regions of heat loading on PFCs (see for example [2, 68]). $<L_{c}>$ is calculated by tracing Monte Carlo particles both forwards and backwards along the field until they intersect a PFC. The connection length of this field line contributes to all grid cells it passes through, and $<L_{c}>$ for a given cell is the arithmetic mean of $L_{c}$ of all unique field lines passing through the cell. The cell in which each magnetic field line is initialized is selected to ensure that each cell contains at least $n_{\text{min}}$ unique trajectories. The location of an initialised field line in a cell is uniformly random. An upper cutoff $L_{c,\text{max}}$ is applied to each field line, since field lines which do not intersect a PFC (intact surfaces around island O-points, for example) would otherwise be traced indefinitely. In the result presented here we select $n_{\text{min}}=100$ and $L_{c,\text{max}}=10^{4}$ meters. The grid resolution used here is 480 points in the radial direction and 1025 points in the poloidal direction. This is several times larger than is typically required for heat deposition calculations but is used here to show the fine structure emerging from the magnetic chaos.
@@
The connection length in the island region is plotted in figure 10 (a), and shows a remarkable pattern of alternating long connection length and short connection length regions. The manifolds of the outer heteroclinic connection are also shown on this plot, and there is a strong correlation between the two. We attribute the small discrepancies between the two do differences in the integrator and coil representation.
@@
The turnstile mechanism helps to explain how this pattern comes about: the region that is not in the turnstile, but is close to the island, will not be mapped out of the island resonance zone in the first many mappings, and has a large connection length. One needs to be in a turnstile to have a low $<L_{c}>$, but since it measures PFC to PFC connection length, only being in one turnstile is not sufficient. We therefore see a lattice of alternating long and short connection length, caused by trajectories that are in both turnstiles alternating with trajectories that are only in one, in an ever repeating pattern that Poincaré dubbed a ’trellis’ [27], made visible in our analysis of the stellarator divertor!
Just like the QUASR configurations, we again see that the turnstile area on the outer heteroclinic connection is much larger than the turnstile on the inner heteroclinic connection. The oscillations of the inner manifold are not visible at all at this scale of magnification. Heat and particles that cross the inner separatrix is channeled along the stable (unstable) manifold, where the above-mentioned trellis structure determines which trajectories eventually intersect the wall. The chaos originating from the turnstile on this outer heteroclinic connection is interesting to study, and is likely to affect the locations where the plasma intersects the wall, but the rate of transport into this region due to chaos in the field is very low due to the small inner turnstile area.
@@
We have seen in all the stellarator configurations, that on each resonance the turnstile on the outer heteroclinic connection is significantly larger than the turnstile on the inner heteroclinic connection. This is likely because the perturbations that induce the chaos are in part due to ripple effects caused by the finite number of discrete coils. The field lines of the outer turnstile pass much closer to these coils, and are more strongly affected by this ripple.
@@
In terms of stellarator divertor design, it would be interesting to understand if this feature is unavoidable, or just a feature of the configurations analyzed. In other words, would it be possible to generate a field with coils, where the turnstile on the inner separatrix is large, potentially larger than that of the outer separatrix? This would enhance the transport from the plasma into the resonance zone at a controllable rate, and would be interesting to contrast with configurations like GYM00+1750 where the chaos is mostly caused by dynamics originating on the outer manifolds.